# Features of the Partition Function of a $\Lambda>0$ Universe

### v1

@@
Abstract
@@ 1 Introduction
The one-loop correction about the four-sphere saddle has been studied in Christensen:1979iy ; Volkov:2000ih ; Anninos:2020hfj ; Law:2020cpj . It was noted in hawking2010general ; Polchinski:1988ua , among other places, that due to the unboundedness of the conformal mode, a phase will generally appear in the resulting one-loop contribution. For the $(d+1)$-dimensional sphere, the one-loop phase goes as (±$(\pm i)\times(\pm i)^{d+1}$, split into a phase ±$\pm i$ associated to an overall constant mode and a phase (±$(\pm i)^{d+1}$ associated to the number of non-isometric conformal Killing vector fields Polchinski:1988ua . The first part of the phase, coming from the constant part of the conformal mode, has been argued to originate microcanonically Anninos:2017hhn ; Anninos:2020hfj ; Maldacena:2024spf , while the second part of the phase has been argued to cancel upon incorporating an observer’s worldline Maldacena:2024spf . [^1] More generally, the phase is modified by the presence of massless higher spin gauge fields, see equation (5.14) of Anninos:2020hfj for a general expression. It was also noted in Anninos:2021ihe that upon coupling a Chern-Simons gauge theory to three-dimensional gravity one can modify and even cancel the overall phase, but this comes at the cost of a parity non-invariant matter sector.
@@
Of the known Einstein metrics with positive curvature, the one with the second largest volume is the Fubini-Study metric on $\mathbb{C}P^{2}$. This spacetime was interpreted as a gravitational instanton in Eguchi:1976db ; Gibbons:1978zy . It is noted in that work, as well as in Hawking:1978ghb , that $\mathbb{C}P^{2}$ does not admit a spin structure. This is due to a non-trivial second Stiefel-Whitney class. Nonetheless, like all orientable four-manifolds hirzebruch1958felder ; teichner1994all , $\mathbb{C}P^{2}$ admits a generalised spin-c structure Hawking:1977ab ; Pope:1980ub . [^2] As such, provided the spectrum of fermions carries judiciously chosen $U(1)$ gauge charges, they can be placed on $\mathbb{C}P^{2}$. For instance, the standard model equipped with sterile right handed neutrinos and a gauged $U(1)_{B-L}$ symmetry can be placed on a manifold with a spin-c structure Garcia-Etxebarria:2018ajm ; Davighi:2022icj . For a discussion of mixed gravitational anomalies associated to the $U(1)_{B-L}$ symmetry, see Putrov:2023jqi . For the mixed gravitational anomaly to be present, one must have that the number of sterile right handed neutrinos differs from the number of generations in the standard model. Further discussion on $U(1)_{B-L}$ can be found in Heeck:2014zfa . In addition, there are several other Einstein metrics including $S^{2}\times S^{2}$, which has been interpreted in terms of the Euclidean continuation of a Nariai black hole Ginsparg:1982rs . It is worth contrasting the multiple saddle-point contributions to $\mathcal{Z}_{\text{grav}}$ with the case of asymptotically flat and anti-de Sitter space in Euclidean signature. In the flat case, the positive action theorem Schon:1979uj ; Gibbons:1979xn indicates that the standard flat metric is the unique asymptotically flat solution to the Euclidean Einstein equations on an $\mathbb{R}^{4}$ manifold. In Euclidean AdS4, upon fixing the asymptotic metric to be conformally equivalent to the flat metric on $\mathbb{R}^{3}$, one generically expects the unique solution to the Einstein equations with $\Lambda<0$ to be Euclidean AdS4, given that the dual CFT3 has a unique conformally invariant vacuum.
@@ 2 Geometric features
[…]
@@
$-\mathcal{S}_{\mathcal{M}_{4}}\equiv\frac{1}{16\pi G}\int d^{4}x\sqrt{g}\left(-R+2\Lambda\right)=-\frac{\Lambda}{8\pi G}V_{\mathcal{M}_{4}}~{}.$ (2.2)
@@
Finally, although we only consider four-dimensions here, it is worth pointing out that in other dimensions the number of known Einstein manifolds is infinite. For instance, in three spacetime dimensions, one has the infinite family of Lens spaces that are given by smooth quotients of $S^{3}$. In five-dimensions, and higher, product metrics of the type $S^{3}\times S^{p}$ are generally present. As such, we automatically have an infinite class of Einstein manifolds. Moreover, in $(2n+3)$-spacetime dimensions, with $n\in\mathbb{Z}^{+}$, one finds increasigly large families of smooth Einstein metrics on $S^{2n+1}$ boyer2005einstein . For instance, there is an infinite family of Böhm metrics on $S^{5}$ bohm1998inhomogeneous enjoying an $SO(3)\times SO(3)$ isometry.
@@ 2.1 Fubini-Study metric on $\mathbb{C}P^{2}$
The volume of the Fubini-Study metric is given by $V_{\mathbb{C}P^{2}}=\tfrac{18\pi^{2}}{\Lambda^{2}}$ which is less than $V_{S^{4}}=\tfrac{24\pi^{2}}{\Lambda^{2}}$, but greater than $V_{S^{2}\times S^{2}}=\tfrac{16\pi^{2}}{\Lambda^{2}}$. It is also known Page:1978vqj that $V_{\mathbb{C}P^{2}\#\overline{\mathbb{C}P^{2}}}\approx\tfrac{150.9}{\Lambda^{2}}$, and that doran2008numerical
@@
Based on the general expression (2.2), the on-shell action of the Fubini-Study metric on $\mathbb{C}P^{2}$ is given by
@@
$\mathcal{S}_{{S^{4}}}>\mathcal{S}_{\mathbb{C}P^{2}}>\mathcal{S}_{{S^{2}\times S^{2}}}>\mathcal{S}_{{\mathbb{C}P^{2}\#\overline{\mathbb{C}P^{2}}}}>\mathcal{S}_{{\mathbb{C}P^{2}\#3\overline{\mathbb{C}P^{2}}}}>...>\mathcal{S}_{{\mathbb{C}P^{2}\#8\overline{\mathbb{C}P^{2}}}}~{}.$ (2.10)
Thus, if $\mathbb{C}P^{2}$ contributes to physical processes, it dominates over $S^{2}\times S^{2}$, and most likely over all other Einstein four-manifolds except $S^{4}$. We are not aware of results on $V_{\mathbb{C}P^{2}\#2\overline{\mathbb{C}P^{2}}}$, but perhaps based on the pattern in (2.10) it will reside somewhere in the interval $[118.4,150.9]$, in units where $\Lambda=1$.
@@ 2.2 Pieces of the partition function of a $\Lambda>0$ Universe
We will verify this explicitly for $\mathbb{C}P^{2}$ below. The overall $i$ in (2.11) comes from the constant part of the conformal mode in the metric fluctuation. The $i^{5}$ from the $S^{4}$ contribution comes from the five conformal Killing vector fields that are not isometries Polchinski:1988ua . The additional $-i$ in ${S^{2}\times S^{2}}$ and ${\mathbb{C}P^{2}\#\overline{\mathbb{C}P^{2}}}$ stems from negative modes in the tensor eigenspectrum Volkov:2000ih ; Young:1983wwy ; Hennigar:2024gbg ; Shi:2025amq ; Ivo:2025yek . It is not known whether the Einstein metric of Chen-LeBrun-Weber chen2008conformally has negative tensor modes, an analysis of various spectra can be found in hall2014spectrum . However, it is known that the Chen-LeBrun-Weber chen2008conformally has only two isometries, and so would contribute the following term to $\mathcal{Z}_{\text{grav}}$
@@
and we have already listed some values of $\chi_{\mathcal{M}_{4}}$. We list them again here for convenience, $\chi_{S^{4}}=2$, $\chi_{\mathbb{C}P^{2}}=3$, $\chi_{S^{2}\times S^{2}}=4$, and $\chi_{\mathbb{C}P^{2}\#k\overline{\mathbb{C}P^{2}}}=3+k$, with $1\leq k\leq 8$ The coupling will add further structure of $\mathcal{Z}_{\text{grav}}$, essentially multiplying each term by a factor $e^{-\vartheta\chi_{\mathcal{M}_{4}}}$. Evidently, to suppress topological contributions (with positive Euler characteristic) one can take $\vartheta\gg 1$. Alternatively, non-spherical topologies can be enhanced upon taking $\vartheta\ll-1$. Similar considerations apply for the Hirzebruch signature,
@@ 2.3 Einstein-Maxwell-$\Lambda$ Theory
where $F=dA$ is the electromagnetic field strength associated to an Abelian $U(1)$ gauge field $A_{\mu}$. This theory can now have additional instantons Brill:1992ce . A simple example is given by the product space $H^{2}\times S^{2}$. This is the Euclidean continuation of the near horizon of a magnetically charged extremal black hole in de Sitter space. The metric and gauge field strength are given by romans1992supersymmetric ; mann1995cosmological ; Booth:1998gf ; Hawking:1995ap
@@
We have selected the root that yields a non-vanishing answer for $A_{2}$ as $\Lambda\to 0$. We also require that $Q^{2}\in(0,\tfrac{1}{4\Lambda})$, such that $A_{1}>0$. We can quotient the hyperbolic space, $H_{2}$, by a freely acting subgroup of $SL(2,\mathbb{R})$ to yield a smooth closed hyperbolic surface, $\Sigma_{h}$, of genus $h$. The resulting Euclidean instanton is then closed, with finite four-volume $V^{(h)}_{Q}$. Minus the Euclidean action is given by
@@
It would be interesting to understand the contribution to the path integral from the sum of all genera. Each Riemann surface will come with a moduli space whose volume is given by the Weil–Petersson metric. [^5] This is reminiscent of computations in two-dimensional quantum gravity, where the resulting sum is asymptotic leading to the inclusion of manifolds with boundaries Polchinski:1994fq ; Ginsparg:1991ws ; Eynard:1992sg ; Shenker:1990uf . Perhaps the same phenomenon occurs for the Gibbons-Hawking path integral, as envisioned in Anninos:2021eit ; Anninos:2024wpy ; Silverstein:2024xnr ; banihashemi2025flat . Other aspects of higher topology in de Sitter space, from a wavefunctional perspective include Anninos:2012ft ; Banerjee:2013mca ; Cotler:2019nbi ; Maldacena:2019cbz ; Anninos:2022ujl ; Aguilar-Gutierrez:2023ril ; Collier:2025lux ; Turiaci:2025xwi . There is also a solution given by $\mathbb{R}^{2}\times S^{2}$, namely
@@
with $Q^{2}=\tfrac{1}{4\Lambda}$, and $A_{2}=\tfrac{1}{2Q^{2}}$. We can quotient the $\mathbb{R}^{2}$ to a torus of complex structure $\tau$. The metric can be viewed as the Euclidean counterpart of the near horizon limit of the ultracold black hole in de Sitter space. The on-shell action now vanishes for all tori! Nonetheless, there is a moduli space of tori that must be integrated over in the path integral, that may lead to mild divergences (reminiscent again of two-dimensional quantum gravity on a torus Bershadsky:1990xb ). This issue, along with the higher genus contributions, will be analysed elsewhere.

### v3

@@
Features of the Partition Function of a $\Lambda>0$ Universe
Dionysios Anninos,1,2 Chiara Baracco,1 Samuel Brian,1 and Frederik Denef3
1Department of Mathematics, King’s College London, Strand, London WC2R 2LS, UK
2Instituut voor Theoretische Fysica, KU Leuven, Celestijnenlaan 200D, B-3001 Leuven, Belgium
3Department of Physics, Columbia University, 538 West 120th Street, New York, New York 10027, U.S.
dionysios.anninos@kcl.ac.uk, samuel.brian@kcl.ac.uk, chiara.baracco@kcl.ac.uk, fmd7@columbia.edu
@@ 1 Introduction
The one-loop correction about the four-sphere saddle has been studied in Christensen:1979iy ; Volkov:2000ih ; Anninos:2020hfj ; Law:2020cpj . It was noted in hawking2010general ; Polchinski:1988ua , among other places, that due to the unboundedness of the conformal mode, a phase will generally appear in the resulting one-loop contribution. For the $(d+1)$-dimensional sphere, the one-loop phase goes as $(\pm i)\times(\pm i)^{d+2}$, split into a phase $\pm i$ associated to an overall constant mode and a phase $(\pm i)^{d+2}$ associated to the number of non-isometric conformal Killing vector fields Polchinski:1988ua . The first part of the phase, coming from the constant part of the conformal mode, has been argued to originate microcanonically Anninos:2017hhn ; Anninos:2020hfj ; Maldacena:2024spf , while the second part of the phase has been argued to cancel upon incorporating an observer’s worldline Maldacena:2024spf . [^1] More generally, the phase is modified by the presence of massless higher spin gauge fields, see equation (5.14) of Anninos:2020hfj for a general expression. It was also noted in Anninos:2021ihe that upon coupling a Chern-Simons gauge theory to three-dimensional gravity one can modify and even cancel the overall phase, but this comes at the cost of a parity non-invariant matter sector.
@@
Of the known Einstein metrics with positive curvature, the one with the second largest volume is the Fubini-Study metric on $\mathbb{C}P^{2}$. This spacetime was interpreted as a gravitational instanton in Eguchi:1976db ; Gibbons:1978zy . It is noted in that work, as well as in Hawking:1978ghb , that $\mathbb{C}P^{2}$ does not admit a spin structure. This is due to a non-trivial second Stiefel-Whitney class. Nonetheless, like all orientable four-manifolds hirzebruch1958felder ; teichner1994all , $\mathbb{C}P^{2}$ admits a generalised spin-c structure Hawking:1977ab ; Pope:1980ub . [^2] As such, provided the spectrum of fermions carries judiciously chosen $U(1)$ gauge charges, they can be placed on $\mathbb{C}P^{2}$. For instance, the standard model equipped with sterile right handed neutrinos and a gauged $U(1)_{B-L}$ symmetry can be placed on a manifold with a spin-c structure Garcia-Etxebarria:2018ajm ; Davighi:2022icj . This remains true for a Higgsed $U(1)_{B-L}$, whereby the low energy effective field theory will also contain Abelian-Higgs vortices Abrikosov:1956sx ; Nielsen:1973cs . For a discussion of mixed gravitational anomalies associated to the $U(1)_{B-L}$ symmetry, see Putrov:2023jqi . For the mixed gravitational anomaly to be present, one must have that the number of sterile right handed neutrinos differs from the number of generations in the standard model. Further discussion on $U(1)_{B-L}$ can be found in Heeck:2014zfa . Potential phenomenological links between $\Lambda>0$ and $U(1)_{B-L}$ are worth further exploration. In addition, there are several other Einstein metrics including $S^{2}\times S^{2}$, which has been interpreted in terms of the Euclidean continuation of a Nariai black hole Ginsparg:1982rs . It is worth contrasting the multiple saddle-point contributions to $\mathcal{Z}_{\text{grav}}$ with the case of asymptotically flat and anti-de Sitter space in Euclidean signature. In the flat case, the positive action theorem Schon:1979uj ; Gibbons:1979xn indicates that the standard flat metric is the unique asymptotically flat solution to the Euclidean Einstein equations on an $\mathbb{R}^{4}$ manifold. In Euclidean AdS4, upon fixing the asymptotic metric to be conformally equivalent to the flat metric on $\mathbb{R}^{3}$, one generically expects the unique solution to the Einstein equations with $\Lambda<0$ to be Euclidean AdS4, given that the dual CFT3 has a unique conformally invariant vacuum.
@@ 2 Geometric features
The solutions are known as Einstein metrics and have constant positive Ricci scalar, $R=4\Lambda$, along with $R_{\mu\nu}=\Lambda g_{\mu\nu}$ by virtue of (2.1). The Einstein metrics reside on a given closed four-manifold, $\mathcal{M}_{4}$. The only known smooth Einstein metric on a topological four-sphere is the round four-sphere, whose volume is given by $V_{S^{4}}=\tfrac{24\pi^{2}}{\Lambda^{2}}$. The other known four-manifolds admitting smooth Einstein metrics are $\mathbb{C}P^{2}$, $S^{2}\times S^{2}$, and $\mathbb{C}P^{2}\#k\overline{\mathbb{C}P^{2}}$, with $1\leq k\leq 8$. [^3] The connected sums $\mathbb{C}P^{2}\#k\overline{\mathbb{C}P^{2}}$ are obtained by removing $k$ balls from $\mathbb{C}P^{2}$ and gluing it to another excised $\mathbb{C}P^{2}$ of opposite orientation at $k$ generic points. For $k=1$ there a is known smooth Einstein metric with constant curvature identified by Page in Page:1978vqj . It describes an $S^{2}$ non-trivially fibered over another $S^{2}$, and admits four Killing vector fields. This metric has smaller volume than the $S^{2}\times S^{2}$ saddle. For $k=2$ an Einstein metric was identified by Chen, Le Brun, and Weber chen2008conformally . For $k\geq 5$ the space of Einstein metrics on $\mathbb{C}P^{2}\#k\overline{\mathbb{C}P^{2}}$ admits a non-trivial moduli space. A survey can be found in anderson2009surveyeinsteinmetrics4manifolds . Of these, only $S^{4}$ and $S^{2}\times S^{2}$ admit a standard spin-structure. However, the remaining four-manifolds admit a spin-c structure Hawking:1977ab . As shown in Garcia-Etxebarria:2018ajm , the spin-c structure implies that the standard model can be placed on this manifold, provided that the BLB-L symmetry, where $B$ is the baryon number and $L$ is the lepton number, is gauged. In what follows, we assume that any matter theory coupled to our gravitational action is compatible with a spin-c structure.
@@
$\mathcal{S}_{\mathcal{M}_{4}}\equiv\frac{1}{16\pi G}\int d^{4}x\sqrt{g}\left(-R+2\Lambda\right)=-\frac{\Lambda}{8\pi G}V_{\mathcal{M}_{4}}~.$ (2.2)
@@
Finally, although we only consider four-dimensions here, it is worth pointing out that in other dimensions the number of known Einstein manifolds is infinite. For instance, in three spacetime dimensions, one has the infinite family of Lens spaces that are given by smooth quotients of $S^{3}$. The sum over such three-dimensional topologies was examined in Carlip:1992wg ; Guadagnini:1994ahx ; Castro:2011xb . In five-dimensions, and higher, product metrics of the type $S^{3}\times S^{p}$ are generally present. As such, we automatically have an infinite class of Einstein manifolds. Moreover, in $(2n+3)$-spacetime dimensions, with $n\in\mathbb{Z}^{+}$, one finds increasigly large families of smooth Einstein metrics on $S^{2n+1}$ boyer2005einstein . For instance, there is an infinite family of Böhm metrics on $S^{5}$ bohm1998inhomogeneous enjoying an $SO(3)\times SO(3)$ isometry.
@@ 2.1 Fubini-Study metric on $\mathbb{C}P^{2}$
The volume of the Fubini-Study metric is given by $V_{\mathbb{C}P^{2}}=\tfrac{18\pi^{2}}{\Lambda^{2}}$ which is less than $V_{S^{4}}=\tfrac{24\pi^{2}}{\Lambda^{2}}$, but greater than $V_{S^{2}\times S^{2}}=\tfrac{16\pi^{2}}{\Lambda^{2}}$. It is also known Page:1978vqj that $V_{\mathbb{C}P^{2}\#\overline{\mathbb{C}P^{2}}}\approx\tfrac{150.9}{\Lambda^{2}}$, that $V_{\mathbb{C}P^{2}\#2\overline{\mathbb{C}P^{2}}}\approx\tfrac{132.8}{\Lambda^{2}}$ hall2014spectrum , and that doran2008numerical
@@
The above equation follows from the fact that the Kähler form of a Kähler-Einstein metric is simply related to the Ricci form and consequently the first Chern class. Based on the general expression (2.2), the on-shell action of the Fubini-Study metric on $\mathbb{C}P^{2}$ is given by
@@
$\mathcal{S}_{{S^{4}}}>\mathcal{S}_{\mathbb{C}P^{2}}>\mathcal{S}_{{S^{2}\times S^{2}}}>\mathcal{S}_{{\mathbb{C}P^{2}\#\overline{\mathbb{C}P^{2}}}}>\mathcal{S}_{{\mathbb{C}P^{2}\#2\overline{\mathbb{C}P^{2}}}}>...>\mathcal{S}_{{\mathbb{C}P^{2}\#8\overline{\mathbb{C}P^{2}}}}~.$ (2.10)
Thus, if $\mathbb{C}P^{2}$ contributes to physical processes, it dominates over $S^{2}\times S^{2}$, and most likely over all other Einstein four-manifolds except $S^{4}$.
@@ 2.2 Pieces of the partition function of a $\Lambda>0$ Universe
We will verify this explicitly for $\mathbb{C}P^{2}$ below. The overall $i$ in (2.11) comes from the constant part of the conformal mode in the metric fluctuation. The $i^{5}$ from the $S^{4}$ contribution comes from the five non-isometric conformal Killing vector fields Polchinski:1988ua . The additional i-i in $S^{2}\times S^{2}$} and $\mathbb{C}P^{2}\#\overline{\mathbb{C}P^{2}}$} stems from negative modes in the tensor eigenspectrum Volkov:2000ih ; Young:1983wwy ; Hennigar:2024gbg ; Shi:2025amq ; Ivo:2025yek . It is not known whether the Einstein metric of Chen-LeBrun-Weber chen2008conformally has negative tensor modes, an analysis of various spectra can be found in hall2014spectrum . [^5] It would be interesting to clarify the appearance of such phases beyond one-loop (or their potential one-loop exactness), as analysed, for instance, in Anninos:2021ene . However, it is known that the Chen-LeBrun-Weber chen2008conformally has only two isometries, and so would contribute the following term to $\mathcal{Z}_{\text{grav}}$
@@
and we have already listed some values of $\chi_{\mathcal{M}_{4}}$. We list them again here for convenience, $\chi_{S^{4}}=2$, $\chi_{\mathbb{C}P^{2}}=3$, $\chi_{S^{2}\times S^{2}}=4$, and $\chi_{\mathbb{C}P^{2}\#k\overline{\mathbb{C}P^{2}}}=3+k$, with $1\leq k\leq 8$ . The coupling will add further structure of $\mathcal{Z}_{\text{grav}}$, essentially multiplying each term by a factor $e^{-\vartheta\chi_{\mathcal{M}_{4}}}$. Evidently, to suppress topological contributions (with positive Euler characteristic) one can take $\vartheta\gg 1$. Alternatively, non-spherical topologies can be enhanced upon taking $\vartheta\ll-1$. Similar considerations apply for the Hirzebruch signature,
@@ 2.3 Einstein-Maxwell-$\Lambda$ Theory
where $F=dA$ is the electromagnetic field strength associated to an Abelian $U(1)$ gauge field $A_{\mu}$. This theory can now have additional instantons Brill:1992ce . A simple example is given by the product space $H^{2}\times S^{2}$. This is the Euclidean continuation of the near horizon of a electromagnetically charged extremal black hole in de Sitter space. For the case of a purely magnetic charge, the metric and gauge field strength are given by romans1992supersymmetric ; mann1995cosmological ; Booth:1998gf ; Hawking:1995ap
@@
We have selected the root that yields a non-vanishing answer for $A_{2}$ as $\Lambda\to 0$. We also require that $Q^{2}\in(0,\tfrac{1}{4\Lambda})$, such that $A_{1}>0$. The flux quantisation condition is such that $Q=\tfrac{\mathfrak{m}}{e}\sqrt{\pi G}$, with $\mathfrak{m}\in\mathbb{Z}$, and $e$ the $U(1)$ gauge coupling. A similar metric can be obtained for the black hole with both electric and magnetic fluxes, but we relegate a complete study to future work.
We can quotient the hyperbolic space, $H_{2}$, by a freely acting subgroup of $SL(2,\mathbb{R})$ to yield a smooth closed hyperbolic surface, $\Sigma_{h}$, of genus $h$. The resulting Euclidean instanton is then closed, with finite four-volume $V^{(h)}_{Q}$. Minus the Euclidean action is given by
@@
It would be interesting to understand the contribution to the path integral from the sum of all genera. Each Riemann surface will come with a moduli space whose volume is given by the Weil–Petersson metric. [^6] This is reminiscent of computations in two-dimensional quantum gravity, where the resulting sum is asymptotic leading to the inclusion of manifolds with boundaries Polchinski:1994fq ; Ginsparg:1991ws ; Eynard:1992sg ; Shenker:1990uf . Perhaps the same phenomenon occurs for the Gibbons-Hawking path integral, as envisioned in Anninos:2021eit ; Anninos:2024wpy ; Silverstein:2024xnr ; banihashemi2025flat . Other aspects of higher topology in de Sitter space, from a wavefunctional perspective include Anninos:2012ft ; Banerjee:2013mca ; Cotler:2019nbi ; Maldacena:2019cbz ; Anninos:2022ujl ; Aguilar-Gutierrez:2023ril ; Collier:2025lux ; Turiaci:2025xwi ; Blacker:2025zca . There is also a solution given by $\mathbb{R}^{2}\times S^{2}$, namely
@@
with $Q^{2}=\tfrac{1}{4\Lambda}$, and $A_{2}=\tfrac{1}{2Q^{2}}$. We can quotient the $\mathbb{R}^{2}$ to a torus of complex structure $\tau$. The metric can be viewed as the Euclidean counterpart of the near horizon limit of the ultracold black hole in de Sitter space. The on-shell action now vanishes for all tori! Nonetheless, there is a moduli space of tori that must be integrated over in the path integral, that may lead to mild divergences (reminiscent again of two-dimensional quantum gravity on a torus Bershadsky:1990xb ). This issue, for the more general case with both electric and magnetic charge, along with the higher genus contributions, will be analysed elsewhere.